The equation determining the liquid pressure in a tank is z = P/rg where z is the depth, P is the pressure, r is the density and g is the acceleration due to gravity. Show that the two sides of the equation are dimensionally the same.


1
Expert's answer
2021-04-15T10:32:52-0400

Consider the equation:


"z = P\/rg."

Let us show that


"\frac{P}{\rho g}=[\text{m}]."

Remember that pressure is


"P=[\text{N/m}^2]=[\text{kg m/(m s)}^2],\\
\rho=[\text{kg/m}^3],\\
g=[\text{m/s}^2]."

Substitute these values:


"\frac{P}{\rho g}=\frac{[\text{kg}][\text{m}][\text{m}^3][\text{s}]^2}{[\text{m}^2][\text{s}^2][\text{kg}][\text{m}]}=[\text{m}]."